OpenStudy (anonymous):

3x + 8y - 36 = 0 The point (x,3/4) is a solution to the equation shown. What is the value of x?

OpenStudy (anonymous):

To find the value of x you have to put in the point (3/4) into the equation and solve for x :)

OpenStudy (mathwizzard3):

The solution is 10. We can substitute 3/4 for y and solve the equation. 3x + 8y - 36 = 0 3x + 8(3/4) = 36 3x + 6 = 36 3x = 30 x = 10

OpenStudy (mathwizzard3):

So.....BABYShawol184 is wrong :/

OpenStudy (mathwizzard3):

Hope that helps ^-^ btw math wizzards can help in other subjects as well :)